Evaluate the integral "\int_0^2 e^{x^3}x^2dx".


1
Expert's answer
2022-06-02T15:01:02-0400
"\int e^{x^3}x^2dx"

"u=x^3, du=3x^2dx"


"\int e^{x^3}x^2dx=\dfrac{1}{3}\int e^{u}du=\dfrac{1}{3}e^u+C"

"=\dfrac{1}{3}e^{x^3}+C"

"\displaystyle\int_{0}^{2}e^{x^3}x^2dx=[\dfrac{1}{3}e^{x^3}]\begin{matrix}
 2 \\
 0
\end{matrix}=\dfrac{1}{3}e^{8}-\dfrac{1}{3}"